# Yihao Wang1,2,4,∗,♢  Pengxiang Ding2,3,4,∗,†  Lingxiao Li1,4,5  Can Cui2,4 Zirui Ge3,4  Xinyang Tong2,4  Wenxuan Song4,6  Han Zhao2,3,4  Wei Zhao2,4 Pengxu Hou6  Siteng Huang2  Yifan Tang1  Wenhui Wang1  Ru Zhang1,🖂

### v1

@@ 1 Introduction
In the past two years, with significant breakthroughs in multimodal LLMs (Karamcheti et al., 2024; Steiner et al., 2024; Liu et al., 2023b; Li et al., 2025), developing robot systems with general perception, understanding, and behavior capabilities has become a key research direction in artificial intelligence. In particular, the emergence of the Vision-Language-Action (VLA) model offers a new solution for enabling robot operations driven by instructions (Kim et al., 2024; Cui et al., 2025; Zhao et al., 2025b; Kim et al., 2025; Song et al., 2025b; Cen et al., 2025; Zhang et al., 2025b; Shi et al., 2025). Research on VLA primarily focuses on extracting multimodal information and aligning it with the action decision space to generate the high-quality actions (Team et al., 2024; Liu et al., 2025; Zhong et al., 2025; Fan et al., 2025).
Current VLA models typically require large-scale embodied data (e.g., Open X-Embodiment (Collaboration et al., 2024) and DROID (Khazatsky et al., 2024)) to pre-train Multimodal Large Language Models (MLLMs) (Especially, VLMs) for task adaptability (Cheang et al., 2024), which is then passed to the designed Policy network (Bu et al., 2024a; Li et al., 2024b) to decode or generate actions for handling the tasks in the diverse environments (Liu et al., 2023a; Mees et al., 2022).
@@ 2.1 Vision-Language-Action (VLA) Models
Recently, leveraging pre-trained Vision-Language Models (VLMs) (Karamcheti et al., 2024; Steiner et al., 2024; Liu et al., 2023b; Li et al., 2025) to control robots for performing various daily tasks has substantially accelerated research in embodied intelligence. This has emerged as a prominent research focus (Black et al., 2025; Intelligence et al., 2025; Shukor et al., 2025; NVIDIA et al., 2025; Liu et al., 2025; Luo et al., 2025; Cheang et al., 2025; Jiang et al., 2025; Ding et al., 2024; Bu et al., 2024b; Fan et al., 2025; Tong et al., 2025). These models are referred to as the VLA models.
Typically, VLA models require large-scale embodied datasets, such as Open X-Embodiment (Collaboration et al., 2024), for pre-training (Liu et al., 2025; Cheang et al., 2025). This process integrates VLMs with a dedicated Policy network (Song et al., 2025a; Li et al., 2024b), allowing the system to decode or generate action sequences for diverse tasks in an end-to-end manner. Moreover, dual-system VLA architectures (Shentu et al., 2024; Bu et al., 2024a; Cui et al., 2025) have recently garnered attention. These methods typically introduce an intermediate latent token to connect the VLMs and the Policy, using an asynchronous mechanism to enhance coordination between the two systems (Zhang et al., 2024a). This design mitigates latency issues during action generation.
@@ 2.2 Bridging from Perception to Action Space
Earlier studies (Kim et al., 2024; Brohan et al., 2023a; b) attempted to directly align perception and action spaces by discretizing actions into tokens. However, this discretization inevitably introduces inherent loss. Recent studies have shifted their focus toward continuous action spaces (Li et al., 2024a; NVIDIA et al., 2025; Black et al., 2025; Shukor et al., 2025; Kim et al., 2025). Based on the types of perceptual features utilized to bridge to the action space, they can be categorized:
@@
Raw features (refer to vision and language representations) are extracted directly from the VLM. Early methods extract representations from the final-layer VLM, operating under the assumption that it encodes the most task-relevant semantic information (Li et al., 2024a; Zhang et al., 2024a). More recent methods leverage the intermediate-layer features within the VLM (Black et al., 2025). They believe that such representations may retain richer multimodal information, thereby benefiting Policy in tasks that demand fine-grained perception or complex reasoning. For example, some studies use features from a middle layer (NVIDIA et al., 2025), the first-half layers (Shukor et al., 2025), or all intermediate-layer features (Black et al., 2025).
@@ 3.1 Preliminary
To build a solid basis for research, we perform experiments of VLA-Adapter on different-scale backbones. The backbones select the Prismatic VLM trained on Qwen2.5-0.5B (Yang et al., 2024), the Prismatic VLM trained on LLaMA2-7B (Touvron et al., 2023), and OpenVLA-7B (Kim et al., 2024) pre-trained on robotic data. The benefit gained from increasing backbone scale is limited in VLA-Adapter. The results are shown in Table 2 of Section 4.1. Therefore, to ensure efficiency, Qwen2.5-0.5B is our default backbone unless otherwise specified.
@@ 3.2 Which Condition Is Essential for Bridging from VL to A?
To ensure compatibility with existing experimental protocols for representative work (e.g., $\pi_{0}$ (Black et al., 2025)), we let the number of Policy layers be equal to that of VLM. At each layer of Policy, the action latent undergoes cross-attention with conditions and self-attention with itself. This iterative process ultimately yields the action output. Details of the Policy can be seen in Section 3.3.
@@ 3.3 Policy with Bridge Attention
To selectively inject certain $\mathcal{{\cal C}}_{t}^{\mathcal{R}}$ into the action space of the Policy, we introduce a learning parameter Ratio $g$ to modulate the influence of $\text{CA}_{1}\left(\widetilde{\bf{A}}^{\tau}_{t},\sigma_{1}(\mathcal{{\cal C}}_{t}^{\mathcal{R}})\right)$. $g$ is initialized to 0 value, and the $\tanh$ activation function is utilized $\tanh(g)\in[-1,1]$ to prevent extreme values from destabilizing the distribution (Zhang et al., 2024b). And then, the three attentions are concatenated to obtain $\widehat{\bf{A}}_{t}^{\tau}$:
@@ 4.1 Necessity of VLA-Adapter
To validate the effectiveness, we compare three kinds of backbones: · B1: The Prismatic VLM (Karamcheti et al., 2024) trained on Qwen2.5-0.5B (Yang et al., 2024). · B2: The Prismatic VLM trained on LLaMA2-7B (Touvron et al., 2023). The first two are different-scale backbones without pre-training on robotic data. · B3: The OpenVLA-7B (Kim et al., 2024) pre-trained on robotic data. We adopted the OpenVLA-OFT bridging paradigm (Kim et al., 2025) for comparison. It is the existing state-of-the-art level method on major benchmarks, including LIBERO-Long (Liu et al., 2023a). The comparison results are shown in Table 2.
Table 2: Effectiveness comparison with OpenVLA-OFT (Kim et al., 2025) on the LIBERO-Long (Liu et al., 2023a). “Fine-tuned” is by LoRA fine-tuning (Hu et al., 2022). Bold represents the best performance. Please note, comparison with the bridge paradigms of $\pi_{0}$ (Black et al., 2025) and GR00T N1 (NVIDIA et al., 2025) has been included in Section 3, so we will not compare it here.
@@ 4.2 Overall Performance on Various Tasks
We selected recently published, comprehensive, and high-performance VLA works as comparison baselines. They are Large: 1. FlowVLA (Zhong et al., 2025), 2. UnifiedVLA (Wang et al., 2025), 3. OpenVLA (Kim et al., 2024), 4. OpenVLA-OFT (Kim et al., 2025), 5. UniVLA (Bu et al., 2025), 6. CoT-VLA (Zhao et al., 2025a), 7. WorldVLA (Cen et al., 2025), 8. TraceVLA (Zheng et al., 2025), 9. MolmoAct (Lee et al., 2025), 10. ThinkAct (Huang et al., 2025), and 11. PD-VLA (Song et al., 2025b); Small: 12. 4D-VLA (Zhang et al., 2025a), 13. SpatialVLA (Qu et al., 2025), 14. $\pi_{0}$ (Black et al., 2025), 15. $\pi_{0}$-FAST (Pertsch et al., 2025), 16. NORA (Hung et al., 2025), 17. SmolVLA (Shukor et al., 2025), 18. GR00T N1 (NVIDIA et al., 2025), and 19. GraspVLA (Deng et al., 2025); Tiny: 20. Seer (Tian et al., 2025), 21. VLA-OS (Gao et al., 2025), and 22. Diffusion Policy (Chi et al., 2023). Their performances are all derived from original references or the reproduction of other published works, ensuring objectivity and accuracy.
@@ Results.
Table 5: Comparison on the LIBERO benchmark. Bold is the best performance. Italics* is suboptimal performance. $\dagger$} represents that the non-based-VLM baselines. “Scratch” is the work without pre-training on robotic data. “Params” is the backbone scale, and its unit is “Billion”. Furthermore, we give the performance of VLA-Adapter on the subtasks. It is shown in Table D1 of Appendix D.
@@
TraceVLA (Zheng et al., 2025) (ArXiv) 7 84.6 85.2 75.1 54.1 74.8
@@
VLA-Adapter (Ours) 0.5 97.8 99.2 97.2* 95.0 97.3
@@
RoboFlamingo (Li et al., 2024b) (ICLR) 3 82.4 61.9 46.6 33.1 23.5 2.48
VPP (Hu et al., 2025) (ICML) 1.5 95.7 91.2 86.3* 81.0* 75.0* 4.33*
@@
Tiny SeerLarge (Tian et al., 2025) (ICLR) 0.57 96.3 91.6* 86.1 80.3 74.0 4.28
@@
VLA-Adapter (Ours) 0.5 99.1 94.6 88.8 82.8 76.5 4.42

### v2

@@ 1 Introduction
In the past two years, with significant breakthroughs in multimodal LLMs (Karamcheti et al., 2024; Steiner et al., 2024; Liu et al., 2023b; Li et al., 2025b), developing robot systems with general perception, understanding, and behavior capabilities has become a key research direction in artificial intelligence. In particular, the emergence of the Vision-Language-Action (VLA) model offers a new solution for enabling robot operations driven by instructions (Kim et al., 2024; Cui et al., 2025; Kim et al., 2025; Song et al., 2025b; Cen et al., 2025; Zhang et al., 2025b; Shi et al., 2025). Research on VLA primarily focuses on extracting multimodal information and aligning it with the action space to generate the high-quality actions (Team et al., 2024; Liu et al., 2024b; Zhong et al., 2025; Fan et al., 2025).
Current VLA models typically require large-scale embodied data (e.g., Open X-Embodiment (Collaboration et al., 2024) and DROID (Khazatsky et al., 2024)) to pre-train Multimodal Large Language Models (MLLMs) (Especially, VLMs) for task adaptability (Cheang et al., 2024), which is then passed to the designed Policy network (Bu et al., 2024a; Li et al., 2024) to decode or generate actions for handling the tasks in the diverse environments (Liu et al., 2023a; Mees et al., 2022).
@@ 2.1 Vision-Language-Action (VLA) Models
Recently, leveraging pre-trained Vision-Language Models (VLMs) (Karamcheti et al., 2024; Steiner et al., 2024; Liu et al., 2023b; Li et al., 2025b) to control robots for performing various daily tasks has substantially accelerated research in embodied intelligence. This has emerged as a prominent research focus (Black et al., 2025b; a; Shukor et al., 2025; NVIDIA et al., 2025; Liu et al., 2024b; Luo et al., 2025; Cheang et al., 2025; Jiang et al., 2025; Ding et al., 2024; Bu et al., 2024b; Fan et al., 2025; Tong et al., 2024). These models are referred to as the VLA models.
Typically, VLA models require large-scale embodied datasets, such as Open X-Embodiment (Collaboration et al., 2024), for pre-training (Liu et al., 2024b; Cheang et al., 2025). This process integrates VLMs with a dedicated Policy network (Song et al., 2025a; Li et al., 2024), allowing the system to decode or generate action sequences for diverse tasks in an end-to-end manner. Moreover, dual-system VLA architectures (Shentu et al., 2024; Bu et al., 2024a; Cui et al., 2025) have recently garnered attention. These methods typically introduce an intermediate latent token to connect the VLMs and the Policy, using an asynchronous mechanism to enhance coordination between the two systems (Zhang et al., 2024). This design mitigates latency issues during action generation.
@@ 2.2 Bridging from Perception to Action Space
Earlier studies (Kim et al., 2024; Brohan et al., 2023a; b) attempted to directly align perception and action spaces by discretizing actions into tokens. However, this discretization inevitably introduces inherent loss. Recent studies have shifted their focus toward continuous action spaces (Liu et al., 2024a; NVIDIA et al., 2025; Black et al., 2025b; Shukor et al., 2025; Kim et al., 2025). Based on the types of perceptual features utilized to bridge to the action space, they can be categorized:
@@
Raw features (refer to vision and language representations) are extracted directly from the VLM. Early methods extract representations from the final-layer VLM, operating under the assumption that it encodes the most task-relevant semantic information (Liu et al., 2024a; Zhang et al., 2024). More recent methods leverage the intermediate-layer features within the VLM (Black et al., 2025b). They believe that such representations may retain richer multimodal information, thereby benefiting Policy in tasks that demand fine-grained perception or complex reasoning. For example, some studies use features from a middle layer (NVIDIA et al., 2025), the first-half layers (Shukor et al., 2025), or all intermediate-layer features (Black et al., 2025b).
@@ 3.1 Preliminary
To build a solid basis for research, we perform experiments of VLA-Adapter on different-scale backbones. The backbones select the Prismatic VLM trained on Qwen2.5-0.5B (Team, 2024), the Prismatic VLM trained on LLaMA2-7B (Touvron et al., 2023), and OpenVLA-7B (Kim et al., 2024) pre-trained on robotic data. The benefit gained from increasing backbone scale is limited in VLA-Adapter. The results are shown in Table 2 of Section 4.1. Therefore, to ensure efficiency, Qwen2.5-0.5B is our default backbone unless otherwise specified.
@@ 3.2 Which Condition Is Essential for Bridging from VL to A?
To ensure compatibility with existing experimental protocols for representative work (e.g., $\pi_{0}$ (Black et al., 2025b)), we let the number of Policy layers be equal to that of VLM. At each layer of Policy, the action latent undergoes cross-attention with conditions and self-attention with itself. This iterative process ultimately yields the action output. Details of the Policy can be seen in Section 3.3.
@@ 3.3 Policy with Bridge Attention
To selectively inject certain $\mathcal{{\cal C}}_{t}^{\mathcal{R}}$ into the action space of the Policy, we introduce a learning parameter Ratio $g$ to modulate the influence of $\text{CA}_{1}\left(\widetilde{\bf{A}}^{\tau}_{t},\sigma_{1}(\mathcal{{\cal C}}_{t}^{\mathcal{R}})\right)$. $g$ is initialized to 0 value, and the $\tanh$ activation function is utilized $\tanh(g)\in[-1,1]$ to prevent extreme values from destabilizing the distribution (Zhang et al., 2023). And then, the three attentions are concatenated to obtain $\widehat{\bf{A}}_{t}^{\tau}$:
@@ 4.1 Necessity of VLA-Adapter
To validate the effectiveness, we compare three kinds of backbones: · B1: The Prismatic VLM (Karamcheti et al., 2024) trained on Qwen2.5-0.5B (Team, 2024). · B2: The Prismatic VLM trained on LLaMA2-7B (Touvron et al., 2023). The first two are different-scale backbones without pre-training on robotic data. · B3: The OpenVLA-7B (Kim et al., 2024) pre-trained on robotic data. We adopted the OpenVLA-OFT bridging paradigm (Kim et al., 2025) for comparison. It is the existing state-of-the-art level method on major benchmarks, including LIBERO-Long (Liu et al., 2023a). The comparison results are shown in Table 2.
Table 2: Effectiveness comparison with OpenVLA-OFT (Kim et al., 2025) on the LIBERO-Long (Liu et al., 2023a). “Fine-tuned” is by LoRA fine-tuning (Hu et al., 2022). Bold represents the best performance. Please note, comparison with the bridge paradigms of $\pi_{0}$ (Black et al., 2025b) and GR00T N1 (NVIDIA et al., 2025) has been included in Section 3, so we will not compare it here.
@@ 4.2 Overall Performance on Various Tasks
We selected recently published, comprehensive, and high-performance VLA works as comparison baselines. They are Large: 1. FlowVLA (Zhong et al., 2025), 2. UnifiedVLA (Li et al., 2025a), 3. OpenVLA (Kim et al., 2024), 4. OpenVLA-OFT (Kim et al., 2025), 5. UniVLA (Bu et al., 2025), 6. CoT-VLA (Zhao et al., 2025a), 7. WorldVLA (Cen et al., 2025), 8. TraceVLA (Zheng et al., 2024), 9. MolmoAct (Lee et al., 2025), 10. ThinkAct (Huang et al., 2025), and 11. PD-VLA (Song et al., 2025b); Small: 12. 4D-VLA (Zhang et al., 2025a), 13. SpatialVLA (Qu et al., 2025), 14. $\pi_{0}$ (Black et al., 2025b), 15. $\pi_{0}$-FAST (Pertsch et al., 2025), 16. NORA (Hung et al., 2025), 17. SmolVLA (Shukor et al., 2025), 18. GR00T N1 (NVIDIA et al., 2025), and 19. GraspVLA (Deng et al., 2025); Tiny: 20. Seer (Tian et al., 2025), 21. VLA-OS (Gao et al., 2025), and 22. Diffusion Policy (Chi et al., 2023). Their performances are all derived from original references or the reproduction of other published works, ensuring objectivity and accuracy.
@@ Results.
Table 5: Comparison on the LIBERO benchmark. Bold* is the best performance, Bold is the suboptimal performance, and Italics is the third best performance. $\dagger$} represents that the non-based-VLM baselines. “Scratch” is the work without pre-training on robotic data. “Params” is the backbone scale, and its unit is “Billion”. We give the performance on subtasks. It is shown in Table D1 of Appendix D. Recently, we have updated the VLA-Adapter-Pro model. Its Policy architecture is the same as Figure 5, and we optimized the implementation. For its details, please see Appendix I.
@@
TraceVLA (Zheng et al., 2024) (ArXiv) 7 84.6 85.2 75.1 54.1 74.8
@@
VLA-Adapter (Ours) 0.5 97.8 99.2 97.2 95.0 97.3
VLA-Adapter-Pro (Ours) 0.5 99.6* 99.6* 98.2* 96.4* 98.5*
@@
RoboFlamingo (Li et al., 2024) (ICLR) 3 82.4 61.9 46.6 33.1 23.5 2.48
VPP (Hu et al., 2025) (ICML) 1.5 95.7 91.2 86.3 81.0 75.0 4.33
@@
Tiny SeerLarge (Tian et al., 2025) (ICLR) 0.57 96.3 91.6 86.1 80.3 74.0 4.28
@@
VLA-Adapter (Ours) 0.5 99.1* 94.6 88.8 82.8 76.5 4.42
VLA-Adapter-Pro (Ours) 0.5 98.5 95.0* 90.5* 85.3* 80.0* 4.50*
@@ Appendix H Effectiveness Analysis of Frozen Backbone
Appendix I Design Details and Performance of VLA-Adapter-Pro
To achieve extreme lightweightness in VLA-Adapter, we shared the projection layer, using the same projection layer for all three attention matrices. This resulted in a low parameter count of 97MB.
In the VLA-Adapter-Pro version, we separated the projection layers for the three attention matrices, allowing different channels to learn differentiated representations. In this case, the parameter count was 207MB. Furthermore, VLA-Adapter-Pro adds Rotary Position Embedding to QK, making the cross-attention more sensitive to position information and more suitable for action generation.
Next, we give the key architecture codes of VLA-Adapter-Pro, as shown below:
class MLPResNetBlock_Pro(nn.Module):
    """One MLP ResNet block with
    separate projections for self, adapter, task
    RoPE, now without FiLM modulation."""
    def __init__(self, dim, num_heads=8):
        super().__init__()
        self.dim = dim
        self.num_heads = num_heads
        self.head_dim = dim // num_heads
        self.ffn = nn.Sequential(
            $n$.LayerNorm(dim),
            nn.Linear(dim, dim),
            $n$.ReLU(),
        # Q (from x only)
        self.q_proj = nn.Linear(dim, dim)
        # Self-Attention: K, V
        self.k_self = nn.Linear(dim, dim)
        self.v_self = nn.Linear(dim, dim)
        # Adapter cross-attention: K, V
        self.k_adapter = nn.Linear(dim, dim)
        self.v_adapter = nn.Linear(dim, dim)
        # Task cross-attention: K, V
        self.k_task = nn.Linear(dim, dim)
        self.v_task = nn.Linear(dim, dim)
        self.o_proj = nn.Linear(dim, dim)
        # gating
        self.gating_factor = nn.Parameter(torch.zeros(1))
        # RoPE
        self.rope = RotaryPositionEmbedding(self.head_dim)
        # ---- FiLM ----
        # # FiLM is useless; to avoid conflict with chkpt, it can be kept as is for now.
        self.film_gen = nn.Sequential(
            nn.Linear(dim, dim * 2),
    def apply_film(self, x, gamma, beta):
        """FiLM: per-channel modulation"""
        return gamma.unsqueeze(1) * x + beta.unsqueeze(1)
    def forward(self, x, h_a=None, h_t=None, p=None):
        h_a: adapter tokens
        h_t: task tokens
        p:   possible conditioning vector (for FiLM)
        g = self.gating_factor
        ratio_g = torch.tanh(g)
        # concat h_a and p
        h_adapter = torch.cat((h_a, p),dim=1)
        h_task = h_t
        B, T, C = x.shape
        K_a = h_adapter.size(1) if h_a is not None else 0
        K_t = h_task.size(1) if h_task is not None else 0
        # Q
        q_1 = self.q_proj(x)
        # self tokens
        k_tokens = self.k_self(x)
        v_tokens = self.v_self(x)
        # adapter tokens
        k_adapter = self.k_adapter(h_adapter)
        v_adapter = self.v_adapter(h_adapter)
        # task tokens
        k_task = self.k_task(h_task)
        v_task = self.v_task(h_task)
        # reshape -> multi-head
        def reshape_heads(t, B, L):
            return t.view(B, L, self.num_heads, self.head_dim).transpose(1, 2)
        q_1 = reshape_heads(q_1, B, T)
        k_tokens, v_tokens = reshape_heads(k_tokens, B, T), reshape_heads(v_tokens, B, T)
        k_adapter, v_adapter = reshape_heads(k_adapter, B, K_a), reshape_heads(v_adapter, B, K_a)
        k_task, v_task = reshape_heads(k_task, B, K_t), reshape_heads(v_task, B, K_t)
        # RoPE
        cos_main, sin_main = self.rope(seq_len=T, device=x.device, dtype=x.dtype)
        q_1, k_tokens = apply_rope(q_1, k_tokens, cos_main, sin_main)
        cos_a, sin_a = self.rope(seq_len=K_a, device=x.device, dtype=x.dtype)
        _, k_adapter = apply_rope(k_adapter, k_adapter, cos_a, sin_a)
        cos_t, sin_t = self.rope(seq_len=K_t, device=x.device, dtype=x.dtype)
        _, k_task = apply_rope(k_task, k_task, cos_t, sin_t)
        # attention scores
        attn_scores = [torch.matmul(q_1, k_tokens.transpose(-2, -1))]
        attn_scores.append(torch.matmul(q_1, k_adapter.transpose(-2, -1)))
        attn_scores.append(torch.matmul(q_1, k_task.transpose(-2, -1)) * ratio_g)
        attn_scores = torch.cat(attn_scores, dim=-1) / math.sqrt(self.head_dim)
        attn_weights = torch.softmax(attn_scores, dim=-1)
        # combine V
        v_list = [v_tokens,v_adapter,v_task]
        v_combined = torch.cat(v_list, dim=2)
        output = torch.matmul(attn_weights, v_combined)
        output = output.transpose(1, 2).contiguous().view(B, T, C)
        output = self.o_proj(output)
        # # ---- FiLM ----
        # gamma_beta = self.film_gen(p)  # [B, 2C]
        # gamma, beta = gamma_beta.chunk(2, dim=-1)  # [B, C], [B, C]
        # output = self.apply_film(output, gamma, beta)
        # residual + FFN
        x = self.ffn(output + x)
        return x
We also present the performance comparison between VLA-Adapter and VLA-Adapter-Pro on 40 subtasks on LIBERO, as shown in Table I1.
Table I1: The specific performance of VLA-Adapter and VLA-Adapter-Pro on the 40 subtasks of four LIBERO (Liu et al., 2023a) suites.
Spatial 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
VLA-Adapter 98.0 100.0 100.0 90.0 96.0 100.0 100.0 100.0 98.0 96.0 97.8
VLA-Adapter-Pro 100.0 98.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.0 99.6
Object 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
VLA-Adapter 98.0 98.0 100.0 100.0 98.0 100.0 98.0 100.0 100.0 100.0 99.2
VLA-Adapter-Pro 100.0 100.0 100.0 100.0 98.0 100.0 98.0 100.0 100.0 100.0 99.6
Goal 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
VLA-Adapter 92.0 100.0 98.0 96.0 100.0 98.0 94.0 100.0 98.0 96.0 97.2
VLA-Adapter-Pro 98.0 100.0 94.0 96.0 100.0 98.0 96.0 100.0 100.0 100.0 98.2
Long 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
VLA-Adapter 96.0 96.0 100.0 98.0 100.0 100.0 84.0 96.0 84.0 96.0 95.0
VLA-Adapter-Pro 92.0 100.0 98.0 96.0 94.0 100.0 94.0 100.0 90.0 100.0 96.4